### v2

# A 2-step Framework for Automated Literary Translation Evaluation: Its Promises and Pitfalls

Abstract
In this work, we propose and evaluate the feasibility of a two-stage pipeline to evaluate literary machine translation, in a fine-grained manner, from English to Korean. The results show that our framework provides fine-grained, interpretable metrics suited for literary translation and obtains a higher correlation with human judgment than traditional machine translation metrics. Nonetheless, it still fails to match inter-human agreement, especially in metrics like Korean Honorifics. We also observe that LLMs tend to favor translations generated by other LLMs, and we highlight the necessity of developing more sophisticated evaluation methods to ensure accurate and culturally sensitive machine translation of literary works.
1 Introduction
Recent studies demonstrate that large language models (LLMs) show strong performance in machine translation Zhang et al. (2023); Jiao et al. (2023). In fact, properties like steerability, context awareness, background knowledge, and context learning capabilities make LLMs suitable for grounded creative tasks like literary translation Karpinska and Iyyer (2023), which is considered the last frontier of machine translation Klemin (2024); Wang et al. (2023a).
While new approaches such as agentic systems are being utilized for literary translation Wu et al. (2024); He et al. (2024), evaluation remains challenging because automated metrics such as BLEU Papineni et al. (2002), BLEURT Sellam et al. (2020) are not capable of capturing the nuance of literary translation Thai et al. (2022). Recently, however, pretrained LLMs have become more common in evaluating machine translation Kocmi and Federmann (2023); Freitag et al. (2023), in lieu of and in collaboration with human evaluation.
Yet, for the agentic translation systems to work at scale, we need an automatic evaluation method first, which will give signal to the LLM agents and pipeline developer whether the translation meets the required criteria or not since human evaluation in each iteration is costly.
Figure 1: The overview of our proposed framework: we evaluate translation of literary works in two stages. In the first step, we focus on semantic and linguistic features of the translation using a human-generated rubric. In the second step, we verify how much of the literary qualities have been captured through the translation by two-agent question answering, where one LLM generates a list of criteria and the other LLM verifies whether those criteria have been satisfied.
In this work, we propose a two-stage framework for literary translation evaluation. In the first stage, we evaluate literary translations with expert-curated Direct Assessment Multidimensional Quality Metrics (MQM) Mariana (2014), while in the second stage, we evaluate the translation using story-specific criteria generated by LLMs for a successful literary translation. A verifier LLM then judges whether the machine translation meets the criteria.
With extensive human evaluation, focusing on English to Korean setting, we demonstrate that while the current LLMs get close to human preference in some dimensions, they still can not replace human evaluation altogether, specifically for some critical dimensions like Korean honorifics. However, we still believe our two-step literary evaluation framework will be useful for future researchers.
2 Background
Literary Translation
Building on James S. Holmes’s 1970s’ foundational theories of literary translation, scholars like Susan Bassett, André LeFevere, and Lawrence Venuti have led the “cultural turn” in translation studies, which emphasizes the importance of cultural and historical contexts in literary translation Snell-Hornby (2006). Literary translation has a special place within broader translation studies for its focus on the aesthetic and affective use of language as well as its engagement with fictional content. Some of the persisting concerns in literary translation include the tension between literal (word-for-word) and free or adaptive (sense-for-sense) Munday et al. (2022), the visibility of the translator Venuti (2017), and the social role of the translator Blakesley (2018). More recently, scholars including Kolb and Macken have examined the potential of the integration of technology and machine translation into literary translation processes Kolb (2023); Macken et al. (2022). Our work presents a pipeline incorporating the cultural and aesthetic aspects of literary translation into the evaluation of machine translation.
Machine Translation Evaluation
String-based metrics like BLEU Papineni et al. (2002) have been widely used in machine translation. Similarly, more recent metrics such as BLEURT Sellam et al. (2020), COMET Rei et al. (2020), and XCOMET Colombo et al. (2023) leverage BERT-style encoder language models Devlin et al. (2019) fine-tuned on translation preference dataset. However, previous work has shown that these metrics are not suitable for literary translation Thai et al. (2022). Departing from earlier trends, Kocmi and Federmann (2023) employed the cutting-edge LLM GPT4 OpenAI (2023) to assess machine translation, where their approach, GEMBA-MQM, set a new benchmark for translation quality estimation at WMT 2023 Freitag et al. (2023). Recent studies from Wu et al. (2024) and Chen et al. (2024) also use GPT4 as a judge for literary translation for web novels and Chinese poetry respectively.
We take a step back and conduct a comprehensive study on the LLMs’ capabilities of fine-grained evaluation for literary translation. A contemporary work, Zhang et al. (2024), has also looked into LLM’s ability to evaluate literary text.
LLM-as-a-Judge
LLM-as-a-judge has emerged as one dominant paradigm in evaluating model outputs Zheng et al. (2023); Liu et al. (2023); Dubois et al. (2023); Chan et al. (2023); Li et al. (2023); Kim et al. (2023). While LLM judges are known to be on par with human judgment in generic tasks in MT-Bench and Chatbot Arena Zheng et al. (2023), several works have highlighted their shortcomings, such as biases in order of the choices Wang et al. (2023b), length of the response Saito et al. (2023), or language of the response Son et al. (2024).
Motivated by earlier research on the effectiveness and limitations of LLM-as-a-judge, we aim to highlight the potential and perils of using LLMs as evaluators within the specific context of literary translation.
Question answering based self-evaluation
Our proposed question-answering-based translation evaluation method relates closely to automatic metrics like TIFA Hu et al. (2023), VPEval Cho et al. (2024), and DSG Cho et al. (2023) for measuring the consistency of text-to-image generation models such as DALL-E Ramesh et al. (2021) and Stable Diffusion Podell et al. (2024). In this line of work, an LLM first generates a list of verification questions from the given user’s prompt, for example “is there a car?" “is the car blue?" and then passes the questions to a vision language model such as BLIP Li et al. (2022) which answers the verification questions. This approach also has a resemblance to how previous works have measured faithfulness in summarization Kim et al. (2024); Wang et al. (2020) and long-form factuality evaluation Min et al. (2023). To the best of our knowledge, our work is the first to explore this question-answering paradigm for machine translation evaluation.
3 Dataset, Metrics & Annotators
Dataset:
We collect 15 short stories translated from English to Korean, representing a diverse range of genres and styles. We then manually align the stories paragraph-wise. The dataset comprises 725 parallel aligned paragraphs across stories, averaging 48.3 paragraphs per story with a standard deviation of 28.9. Each paragraph contains an average of 6.8 sentences, with a standard deviation of 6.9, in English. A full list of the stories and additional details can be found in Table 7 in Appendix A.
Metrics:
We use expert human evaluation to verify the efficacy and viability of our proposed framework. In all steps, we use Kendall’s Tau and Spearman correlation to measure the correlation with human judgment. We also report the mean squared error (MSE) with the scores provided by models with human annotators’ scores.
Human Annotators:
Three literary experts with extensive experience in both English and Korean literature develop and validate the framework presented in Section 4. They also carry out the annotations in Section 5 to validate the framework. All annotators were female and native Korean speakers.
4 Two-Step Literary Translation Evaluation Framework
Even experts continue to disagree on what constitutes a good literary translation. The nuanced interplay of language and culture in literature often produces multiple, equally valid translations Sager (1998). Moreover, translators sometimes have to make the difficult decision to deviate from the original text to culturally adapt the material so that the essence of the work is conveyed. A good literary translation often diverges from a faithful, word-to-word rendering. Evaluating literary translation thus demands unique criteria and standards compared to other machine translation tasks.
In this work, we propose a two-step framework for evaluating literary translation (see Figure 1):
Step 1
RULER: RUbric-based Literary Evaluation Rating is a Likert scale of 1 to 5 with four criteria, which resemble Multidimensional Quality Metric (MQM) Mariana (2014), for evaluating translations based on rubrics.
Step 2
VERSE: A VERification-based Story specific Evaluation that asks story-specific literary questions that consider the distinct characteristics needed in successful translation of literary works.
4.1 Step 1: RULER
We first conduct a small-scale qualitative evaluation of machine-translated literary texts to identify the capabilities that evaluation models should possess, with a particular focus on the types of errors they should detect. For this, we select five stories (marked dark gray in Table 7) and translate them using various large and small language models, as well as commercially available translation services, like Google Translate, to get a diverse set of translations. We observe that LLMs show inconsistent quality in doing literary translation. Larger models are more likely to produce remarkably sophisticated translations, while smaller ones often make unexpected errors, even with relatively simple passages. Our analysis identifies four major types of errors machine translation systems make when translating literary works, some of which are especially pronounced in English-to-Korean translations: word/phrase/expression-level issues, context-related issues, issues relevant to honorifics, and form-related issues (see examples in Table 1). We provide a detailed explanation of these categories with concrete examples in Appendix B.
Table 1: Examples of LLM mistranslations for each category. Examples are from Chesnutt’s “The Passing of Grandison”(1899), a short story on the social construction of racial dynamics.
From this initial qualitative evaluation of translations, we identify four categories that are critical to the foundational quality of English-to-Korean literary translations (listed below). We then develop a fine-grained rubric to systematically assess translations, focusing on these categories designed to detect common mistake patterns in literary translation.
  • Lexical Choice Proper use of idiomatic expressions and general word choices that demonstrate fluency in the target language.
  • Proper Use of Honorifics in Dialogues Selection of verb endings that accurately reflect the status of and relationships among the participants in a conversation, ensuring natural-sounding dialogue in Korean. Note that this applies only to paragraphs containing dialogue or other expressions of register in the description.
  • Syntax and Grammar Grammatical accuracy of sentences including proper translation of embedded clauses, conjunctions, sentence connectors, internal logic of sentences, and punctuations, which ultimately makes the translation read naturally in Korean.
  • Content Accuracy Preservation of the text’s meaning, with particular emphasis on proper and comprehensive translation of literary expressions in the original text. While accurate translation might involve some additions or omissions for clarity, it should avoid altering or omitting substantial words or phrases that are essential to the content and context of the original.
The detailed rubric for each criterion can be found in Appendix C.
4.2 Step 2: VERSE
Eliminating common errors is insufficient for achieving high-quality literary translation. A successful translation must also maintain the essential and unique literary qualities of the original work, which can vary significantly between different texts. To capture such nuances, we propose VERSE, a question-answering-based evaluation framework. We prompt a high-performing LLM (in our case GPT4-o OpenAI (2024)) to generate a list of verification questions that must be satisfied in a good translation. We instruct the question-generating model to focus on the literary aspects of the passage, with a brief summary of the story provided for context [^1] The full prompt is provided in Appendix J.1. Figure 1 provides examples of the questions generated by the LLM222additional examples are available in Appendix K. A thorough examination by native Korean speakers confirmed that LLMs are capable of generating questions that effectively address key considerations in translating literary works with varying styles and other quirks.
Another LLM then evaluates the machine translation based on these questions. The grading was done on a 1 to 3 scale, where 1 indicates the criteria in the question have not been satisfied at all, 2 indicates partial satisfaction, and 3 indicates full satisfaction of the criteria.
Information about the quality of LLM generated question can be found in Appendix H.1.2.
Clustering LLM Generated Questions
To provide better interpretability for the LLM generated questions in VERSE, we classify the questions into nine categories using an LLM. For all our experiments, the questions are classified in a zero shot manner by GPT4-o. The categories [^3] Notice that the categories are not necessarily mutually exclusive and one question might fit into more than one category., as shown in Figure 2, are defined by a human expert after qualitatively examining many evaluation questions. To validate the accuracy of the LLM classification, one human expert manually annotates 150 questions into the nine categories in which GPT4-o scores  77% Top-1 match in accuracy. An examination of the misclassified examples reveals that most misclassifications also fall into a closely related category  [^4] See confusion matrix in Figure 5.
Figure 2: Different stories require different kinds of attention to ensure that the literary essence of the work is preserved in the translation. The categories were found after classifying them in zero shot manner. See Section 4.2. The radial axis represents the percentage of questions belonging to that category for that particular story.
5 Results
In this section, we evaluate the effectiveness of the two-step framework proposed in Section 4 using human annotation.
Human Annotation
We first create a diverse set of translations by machine translating the stories in our datasets with various LLMs and commercial translation service (e.g., Google Translate). The LLMs ranged from small 7B models to GPT4-o and Claude Sonnet 3.5555similar to what we described in Section 4.1. We also vary different prompting techniques (sentence level, paragraph level, providing extra summary, zero shot, few shot (see Appendix G)) to test the generalizability.
We then randomly select 20 ( machine translated paragraph) pairs from each of the 10 stories marked gray in Table 7 for human annotation (200 paragraphs in total). We also randomly select three evaluation questions corresponding to those paragraphs for annotation for Step 2 (600 questions in total).
Because of the subjective nature of evaluating literary translation, we create an extensive annotation guideline alongside the rubric so that annotators have the same understanding of the given task. The annotation guideline and annotation interface can be found in Appendix E. Table 2 shows that the annotators have high inter-annotator agreement. We also ask the annotators to re-annotate a subset of 40 paragraphs after four weeks of the first annotation, and the intra-annotator agreement was  0.7.
Criteria Kendall’s Spear. MSE Krip.
Category $\tau$ $\rho$ $\alpha$
Honorifics .70 .74 .89 .73
Lex. Choices .72 .80 .71 .79
Syntax .66 .75 .88 .74
Content .72 .81 .77 .79
LLM QnA .64 .69 .42 .69
Table 2: Inter-annotator agreement metrics across categories. Higher values indicate better agreement for Kendall’s $\tau$, Spearman’s $\rho$, and Krippendorff’s $\alpha$, while lower values are better for MSE. The values are averaged over three annotators. Notice that the last row, LLM QnA, belonging to Step 2, is measured between 1 to 3 while the others, belonging to Step 1, are measured on a 1 to 5 scale.
Experiment Setting
We compare several open-weight and proprietary models, which are known for their good performance in Korean, as evaluators for our experiments. We set the temperature to 0.0 for all experiments in this section for reproducibility and report the numbers from a single run from here on.
Baseline
According to the common practice of literary machine translation Karpinska and Iyyer (2023); Chen et al. (2024); He et al. (2024) , we report COMET Rei et al. (2022), BERTScore Zhang* et al. (2020) and BLEURT Sellam et al. (2020) as the baseline. However, the comparison may not be entirely appropriate because these metrics are meant to be used in sentence-level evaluation, and our dataset is aligned at paragraph-level.
5.1 RULER: Can LLMs Capture Common Mistakes in Literary Machine Translation?
Crit. Label Sup. H- GPT4-o- H- GPT4-o-
F1 F1 Acc Acc
Honor. 1 31 0.79 0.07 0.81 0.66
2 3 0.15 0.10
3 7 0.11 0.04
4 11 0.22 0.14
5 140 0.91 0.88
Lexical 1 16 0.54 0.43 0.54 0.47
2 32 0.47 0.45
3 45 0.45 0.37
4 47 0.50 0.52
5 52 0.70 0.51
Syntax 1 23 0.61 0.39 0.52 0.43
2 32 0.55 0.45
3 49 0.50 0.27
4 43 0.38 0.44
5 45 0.60 0.56
Content 1 19 0.57 0.19 0.54 0.46
2 32 0.41 0.47
3 47 0.45 0.22
4 48 0.47 0.43
5 46 0.73 0.70
Table 3: Performance comparison across different evaluation criteria. H-F1 represents the average F1-score between human annotator pairs (3 pairs), while LLM-F1 shows the average F1-score between human annotators and GPT4o (3 pairs). H-Acc and LLM-Acc show the average accuracy for human-human and human-LLM comparisons respectively. Support (Sup.) shows the average number of instances for each label across three annotators (rounded).
Honorifics Lexical Syntax Content Aggregated Avg.
Model $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE Avg. $\tau$ Avg. $\rho$ Avg. MSE
gpt-4o .50 .55 2.10 .68 .77 0.71 .62 .71 0.95 .70 .79 0.81 .62 .70 1.14
claude-3-5-sonnet .39 .41 2.30 .58 .65 1.17 .57 .65 1.20 .57 .64 1.38 .53 .59 1.51
claude-3-5-haiku .41 .44 1.86 .54 .60 1.37 .57 .64 1.23 .55 .62 1.42 .52 .58 1.47
gpt-4o-mini .26 .29 2.84 .62 .70 0.93 .60 .68 0.91 .60 .67 1.10 .52 .58 1.44
claude-3-haiku .20 .21 2.24 .45 .50 1.57 .42 .47 1.47 .50 .56 1.55 .39 .43 1.71
gemma-2-27b-it .33 .36 2.16 .46 .51 1.27 .48 .54 1.24 .48 .54 1.26 .44 .49 1.48
gemma-2-9b-it .27 .28 2.44 .42 .47 1.41 .35 .40 1.48 .52 .58 1.35 .39 .43 1.67
llama-3-8b-it -.02 -.02 2.36 .37 .41 1.58 .28 .32 1.67 .21 .23 2.09 .21 .23 1.92
COMET .09 .11 .27 .34 .24 .32 .27 .35 .22 .28
BERTScore .12 .15 .23 .29 .22 .28 .24 .31 .20 .26
BLEURT .13 .16 .22 .28 .20 .26 .22 .29 .19 .25
Table 4: (Step 1) Performance comparison of different models across different criteria. $\tau$: Kendall’s Tau, $\rho$: Spearman’s correlation, MSE: Mean Squared Error. Best results in each column are in bold. Higher is better for $\tau$ and $\rho$ and lower is better for MSE. The exact model versions can be found in the appendix. All values were computed over 200 samples and run to run variations were minimal. Three traditional machine translation baselines are given below.
We first prompt the language model with the human-curated rubrics for each criterion and ask it to give a score. The LLM is also given one gold translation for reference. The full prompt can be found in Appendix D. As shown in Table 4, across all the criteria, LLM-based evaluation metric outperforms traditional metrics, however, even the best evaluator, GPT4-o does not match human-level correlation reported in Table 2. Our findings show that larger proprietary LLMs tend to do better than smaller models.
In particular, a careful analysis reveals that GPT4-o and all other models completely fail to notice serious mistakes in honorifics. Table 3 shows that GPT4-o often gives high scores for translations that received low scores from humans regarding honorifics (notice the low F1 score for labels 1, 2 and 3).
Korean honorifics play a crucial role in dialogue-based literary texts, drastically impacting the reading experience when improperly translated. Our evaluation rubric assigns a default score of 5 to non-dialogue passages, which means a significant number of passages automatically get 5. Thus, in our case, it is important for the models to detect when the honorifics get translated incorrectly. However, as Figure 6 shows GPT4-o’s predicted label distribution does not match well with human-to-human agreement. In contrast, GPT4-o does a better job at matching human preference distribution for Syntax, Content Accuracy, and Lexical choices (Figure 7), however, they still tend to provide overoptimistic estimations, a phenomenon captured by higher MSE compared to inter-human. We explore this phenomenon further in Section 6.
We conduct ablations of each component of the pipeline and reach the following conclusion:
  • CoT does not improve performance and, in fact, sometimes hurts (Table 9).
  • Specific grading rubric is important for some models, and not so much for others (Table 10).
  • Providing a reference translation does not help much and in fact, for some models, it is better not to provide one (Table 10).
  • Providing few shot examples has little to no effect (Table 11).
Appendix H digs into the observations in more detail.
5.2 VERSE: Can LLMs Answer Literary Questions with Subtle Nuance?
Similar to Step 1 RULER, we prompt the language models with human curated rubric and a reference human translation and ask them to give a score.
Model Kendall Spearman MSE
gpt-4o .49 .54 0.61
claude-3-5-sonnet .51 .54 0.73
claude-3-5-haiku .37 .40 0.77
gpt-4o-mini .37 .40 0.61
claude-3-haiku .20 .21 0.82
gemma-2-27b-it .13 .14 0.71
gemma-2-9b-it .09 .09 0.71
Llama-3.1-8B-it .05 .05 1.08
COMET .23 .30
BERTScore .19 .24
BLEURT .19 .24
Table 5: (Step 2) VERSE: Performance comparison across different models on evaluating whether a particular criterion was satisfied in the translation or not. Usually, larger models do better than smaller ones. The results are for zero shot no CoT setting. Baselines are provided below.
As shown in Table 5, language models are much better at correlating with human preference in understanding literary translation than the traditional MT metrics, however, there remains a gap between expert human-level agreement (Table 2). In-depth analysis of GPT4-o (the best-performing model above) performance shows that the model has a particularly lower recall than inter-annotators for Score 1 than for Score 2 and 3 (Table 8), which means GPT4-o cannot effectively spot translations that critically fail to satisfy specific criteria. However, the recall drop was not as significant as in the case of honorifics in Step 1.
We identify the following key components of LLM’s ability to answer fine-grained questions:
  • CoT only helps certain models and hurts a few (Table 12).
  • Human written rubrics (Appendix D) are critical (Table 10) and correlation drops to near zero without them.
  • Providing a reference translation does not help evaluation, in fact not providing is slightly better (Table 13).
  • Providing few shot examples does not make a significant improvement across the models (Table 14).
Appendix H digs more into these findings.
6 Discussion: Promise and Pitfall of the Proposed Framework
While larger language models perform clearly better than traditional machine translation metrics, they still fall behind inter-human correlation. We, therefore, further examine what the use cases and limitations of our framework for automated literary translation evaluation may be.
6.1 Experiment Setting
For all the experiments in this section, we use the GPT4-o as the evaluator model for all experiments in zero-shot no CoT setting. We also did not use reference human translation because, in Step 2, reference-free configuration performs slightly better [^6] We also report the evaluation scores computed with reference translation in the prompt in Appendix H.
Evaluated Translations
To test the efficacy of our framework, we use several language models and Google Translate to translate 10 stories from our dataset. Following the best practices of Karpinska and Iyyer (2023), we prompt language models with five examples and then generate translations. We also append a GPT4o-generated summary of the story for context.
From here on, we use $\mathrm{LM}_{E}$ and $\mathrm{LM}_{G}$ to differentiate between the evaluator language model and translator language models.
6.2 Fine Grained Literary Evaluation
Figure 3: Radar plot for scores (as percentage) across different categories of translation for each model in The Philosopher story. The radial axis begins at 40%. See Figure 8 for more examples.
VERSE (Step 2) of our framework permits a novel fine-grained interpretation into literary translation, which is otherwise hard to capture through a single number or traditional machine translation metric like fluency, adequacy, etc. Figure 3 shows one such use case of such approach to quickly inspect and visualize different dimensions present in the translation. Moreover, because the language model evaluates  10 questions per paragraph, VERSE estimations are also robust.
6.3 LLMs prefers LLM generated translation
Table 6 shows the aggregated scores for each Step 1 criteria for each model (see Table 16 for Step 2). Note that since our proposed framework can be run in reference free fashion, we also evaluate the given reference human translations with our framework.
The results are surprising. Overall, Claude-3.5 Sonnet, as $\mathrm{LM}_{G}$, gets the highest scores, followed by GPT4-o, with almost perfect scores in several dimensions. Human translation receives lower scores than the bigger models. To better understand what is going on, we sample 50 paragraphs from the story “What is Remembered", whose human translation receive a particularly lower score than the best performing $\mathrm{LM}_{G}$ translation by ($\mathrm{LM}_{G}$) Claude Sonnet 3.5 (See Figure 8). In a controlled A/B setting, one human annotator picked the gold translation as the better one 39 out of 50 times, and the Claude Sonnet 3.5 translation was picked only 4 times.
An inherent characteristic of human translation may explain why $\mathrm{LM}_{E}$ gives human translation lower scores. Human translations often deviate significantly from the surface-level linguistic features of their original text in order to properly convey the essence of the text. Human inspection reveals that this is indeed the case for the story examined above. In the cases of other stories, too, human translations often prove to read more naturally than the best machine translations in most cases, even when the best machine translations were sufficiently accurate.
We observe the same for Step 2, where translations by large models are rated higher than human translations.
This suggests that while LLMs may be good at spotting significant differences in translation qualities, they still often fail to distinguish translations where the difference is more subtle and nuanced. This can be observed from the relatively lower scores provided to Google Translate and GPT3.5, which usually produce very poor Korean translations.
Transl. Model Hon. Syn. Lex. Con.
Human 98.97 88.21 89.74 83.08
GPT-4o 99.49 97.44 94.36 97.95
Claude-3.5-Sonnet 100.00 98.97 98.97 98.97
Claude-3.5-Haiku 100.00 94.87 96.41 96.92
Claude-3-Haiku 98.46 88.72 85.64 89.23
GPT-4-Mini 99.49 90.26 88.21 92.31
GPT-3.5 91.79 57.44 56.92 60.00
Google Translate 97.95 71.79 69.23 76.41
Table 6: Average scores (in percentage) across different models on four criteria of translation quality, computed over the 10 evaluation stories (501 paragraphs). Note that while LLMs can discriminate between very good and very bad translations (eg, GPT4o vs GPT3.5), its ability to distinguish between high-quality translations is limited.
Generalizability of Our Framework
While we keep our analysis limited to English to Korean, the overall structure of evaluation can be used for other languages, provided LLMs have a reasonable understanding of the target language. Our framework is simple, easy to implement, and provides interpretability suited specifically for literary translation. Much of the pipeline, especially, VERSE is language agnostic and works without much human supervision. For RULER, we created the rubric with the Korean language in mind, however, RULER can easily be adopted with modification for other languages. While gaps remain in current LLMs’ ability to detect some critical issues in literary translation, we believe this paradigm will be helpful for future researchers and practitioners to evaluate literary translation.
7 Conclusion
In this work, we propose and evaluate the feasibility of using LLM as an automated tool for fine-grained evaluation of literary text and found that current LLMs, while getting closer to human-level judgment than traditional metrics and providing useful interpretable insights, still lag behind human judgments, especially in dimensions like honorifics. The findings call for more research in developing metrics for evaluating long-form translations, especially in the literary domain.
Limitations and Future Work
Our work only tests the framework on one pair of languages, English to Korean. Future works should incorporate other language pairs, which provide different levels of performance. Especially, it remains to be seen how well LLMs can generate literary-specific questions for other non-English languages and then answer them. It is also important to see whether LLMs can take into account different cultural contexts of the source language in the translation. It is possible that LLMs will not be able to evaluate high-quality questions in low-resource languages or languages that they haven’t been trained on extensively. Similarly, it is also possible that the framework works even better in higher-resource languages.
Appendix A Datast detail
Our selected collection of short stories spans nearly two centuries, offering a survey of diverse literary movements and genres in the English language. From classic nineteenth-century gothic and popular suspense stories like Edgar Allan Poe’s “The Masque of the Red Death” (1842) and Charles Dickens’s “The Signal-Man” (1866), to modernist text like Virginia Woolf’s “The Legacy” (1944) and contemporary realist works such as Jhumpa Lahiri’s “Hell-Heaven” (2004) and Andrew Porter’s “Austin” (2017). The genre of speculative fiction tradition is represented by Ursula K. Le Guin’s “The Ones Who Walk Away from Omelas” (1973), Ted Chiang’s “Tower of Babylon” (1990), and Octavia Butler’s “Amnesty” (2003), all of which differ in writing style, while works like Charles Chesnutt’s “The Passing of Grandison” (1899), Katherine Mansfield’s “The Garden Party”(1922) and Tobias Wolff’s "The Deposition" (2006) represent works in which historical and cultural detail contribute greatly to their style. As a collection, these stories cover a wide-range of evolving narrative styles and techniques in English literature.
Paragraphs from five of the fifteen stories (shaded dark gray on top) were used to create the few-shot examples for all the experiments in Section 4 and 6. The other five works (lighter gray) were part of the human evaluation in 4 but their paragraphs were not part of the few-shot examples. We leave the five other stories solely for the purpose of our analysis in Section 6.
The names of the stories and other details can be found in Table 7.
Title Author #P Translator(s) Orig. Trans.
Austin Andrew J. Porter 26 Eun Young Min 2017 2024
The Passing of Grandison Charles W. Chesnutt 56 Ki Wook Han 1899 2010
The Ones Who Walk Away
from Omelas
Ursula K. Le Guin 14 Yong Jun Choi 1973 2014
The Garden Party Katherine Mansfield 44 Young Hee Kim 1922 2010
Real Artists Ken Liu 84 Seong Ju Jang 2011 2024
The Philosopher Sherwood Anderson 29 Seon H. Kim, Young W. Park 1919 16, 19
A Retrieved Reformation O. Henry 36 W. D. Kim, Jeong A Ko 1903 03, 14
Amnesty Octavia Butler 84 Soo Hyun Lee 2003 2016
Tower of Babylon Ted Chiang 107 Sang Hoon Kim 1990 2016
The Red Masque of Death Edgar Allan Poe 14 D. H. Jeon, S. H. Jeon 1842 09, 14
Hell-Heaven Jhumpa Lahiri 34 Sang Mi Park 2004 2009
The Legacy Virginia Woolf 16 Young Hee Kim 1944 2010
The Signal Man Charles Dickens 52 Young Hee Kim 1866 2010
What is Remembered Alice Munro 90 Jeong Eun Seo 2001 2020
The Deposition Tobias Wolff 39 Jae Kyung Lee 2006 2012
Table 7: The details of the stories being used in our work. The number of paragraphs and when they were published are shown. Three stories in our corpus have two different translations, for which we included the names of both translators. Gray row stories were used for human evaluation experiments in Section 5.
Appendix B Details of Common Issues found in Machine Translation
Word/Phrase/Expression Level Issues
This type of error occurs most frequently and significantly impacts overall translation quality. While LLMs excel at literal translations, they struggle with natural word choices in the target language, particularly when dealing with figurative expressions. This often results in the selection of awkward or uncommon verbs in Korean; although the overall content remains comprehensible, these choices critically diminish the quality of translation. Notably, problems often arise with polysemous words or idiomatic expressions, seriously impeding reader comprehension. Additional common errors include inappropriate use of pronouns, such as referring to people as “{CJK*}UTF8mj그것 (it),” reflecting the essential difference in the linguistic structure between English and Korean.
Context Related Issues
This type of error presents more complex and serious challenges arising from cultural and historical differences between the source and target languages. For instance, because terms like “abolitionist” in “The Passing of Grandison” do not have an equivalent historical term in Korean language, LLMs end up offering incomprehensible expressions. Similarly, the expression “border states” often loses its geographical context (mistranslated as “{CJK*}UTF8mj국경주–states close to national borders” in Korean). LLMs also struggle with specialized terminology and historical references, as shown in mistranslations of terms like “bootjack” and “nightcap.”
Form-related issues
This third category includes subcategories concerning punctuation and grammar, as well as structure and formatting. Punctuation and grammar refer to problems with punctuation changes that alter meaning. For example, when LLM translation replaces semicolons with periods, it alters the original text’s meaning. Likewise, changes in the sentence structure in the translated text can lead to incomplete or awkward results. This subcategory includes problems with splitting long sentences unnecessarily, which results in awkward structures when handling complex sentences. Other issues in this category included structure and formatting. These appeared when formatting the dialogue and narrative flows. The incorrect handling of sentence structures in the LLM translation led to awkward sentences, negatively affecting readability. Form-related issues also include transliteration, consistency issues, and omissions. LLMs transliterate Western names inconsistently in the translation, resulting in different name variations throughout the translated text. (i.e., translating “Grandison” as “{CJK*}UTF8mj그라디슨" and “{CJK*}UTF8mj그란디손” within a single translation) LLM translations may eliminate a sentence and expression or add arbitrary content different from the original text. They also leave certain words untranslated, sometimes including various words or phrases in random languages within the translation.
Honorifics
Errors in this category are notable in English-to-Korean LLM translations regarding features of Korean, in which different levels of formality are presented through various verb endings. Mistranslation of honorifics and formality appear most frequently in dialogues, reflecting a misinterpretation of the relationships between characters, or inconsistent use of honorifics within a single character’s speech. Translation issues in the honorifics category critically affect the reading experience because it encompasses naturalness, as well as proper understanding of character power dynamics and subtle psychological contexts.
Appendix C Detailed Rubric for Each Criteria (Step 1)
C.1 Lexical Choices
Proper use of idiomatic expressions and general word choices that demonstrate fluency in the target language
  • Score 5 The translated text chooses words in the target language that best suit the authorial intention of the original text, conveying meaning in a properly idiomatic and fluent manner. The translated text thereby provides readers with an equivalent reading experience and emotional impact when compared with the original text.
  • Score 4 The translated text mostly selects appropriate words that match the authorial intention, maintaining idiomatic expressions and fluency in the target language. The translated text generally preserves the reading experience and emotional effect of the original text.
  • Score 3 The translated text makes some less accurate or awkward word choices that compromise idiomatic fluency while demonstrating good comprehension of the original text. Lexical choices in the translation somewhat negatively affect readers’ experience and emotional engagement with the text.
  • Score 2 The translated text includes multiple inappropriate word choices that compromise idiomatic fluency to the extent that it makes it difficult to understand the authorial intention of the original text. Lexical choices in the translation somewhat negatively affect the reading experience and emotional engagement.
  • Score 1 Translated text makes clearly awkward word choices that reveal a critical lack of understanding of idiomatic and natural expressions in the target language, seriously hindering readers’ comprehension and engagement with the text. Lexical choices in the translation reflect critical miscomprehension of the authorial intent of the original text.
C.2 Proper Use of Honorifics in Dialogues
Selection of verb endings that accurately reflect the status of and relationships among the participants in a conversation, resulting in natural-sounding dialogue in the Korean language.
  • Score 5 The translated text properly captures the relationship between the speaker and the listener(s) of the conversation. The translated text also accurately reflects the culturally and historically specific elements such as titles and modes of appellations. The level of politeness aligns appropriately with the situational context, familiarity, and character.
  • Score 4 The translated text mostly captures the relationship between the speaker and the listener(s) of the conversation, with minor inaccuracies in culturally and historically specific elements such as titles and modes of appellations. The level of politeness mostly aligns with the situational context and social status of characters. There may be minor inconsistencies but the dialogue accurately represents the characters’ familiarity and relationships.
  • Score 3 The translated text somewhat captures the relationship between the speaker and the listener(s). But the translated text includes noticeable errors in culturally and historically specific elements such as titles and modes of appellations. The level of politeness partially aligns with the situational context and the social status of or relationships of the characters. There are some inconsistencies or awkwardness in the dialogue that are noticeable for native speakers of Korean.
  • Score 2 The translated text barely captures the relationship between the speaker and the listener(s). The translated text demonstrates significant misunderstandings of culturally and historically specific elements such as titles and modes of appellations. The level of politeness barely aligns with the situational context, leading to a misrepresentation of the characters’ relationships and awkwardness in the target language.
  • Score 1 The translated text completely misunderstands the relationship between the speaker and the listener(s) of the conversation and fails to reflect cultural and historical specificity of the context. The level of politeness in the used language fails to align with the situational context. Inconsistencies or misinterpretations of the appropriate degree of formality significantly undermine the dialogue’s naturalness in Korean.
C.3 Syntax and Grammar
Grammatical accuracy of sentences including proper translation of embedded clauses, conjunctions, sentence connectors, internal logic of sentences, and punctuations, which ultimately makes the translation read naturally in the target language.
  • Score 5 The translated text accurately conveys the meaning, logic, and nuances of the original text. Embedded clauses, conjunctions, and sentence connectors are translated meticulously so that the translation reads naturally in the target language without awkward or confusing sentence structures. The number of sentences in the translated text may vary from the number of sentences in the original text if it enhances the naturalness and readability of the translation.
  • Score 4 The translated text generally preserves the meaning, logic, and nuances of the original text. The translation of embedded clauses, conjunctions, and sentence connectors reads naturally, with only minor instances of awkwardness or ambiguity. Variations in the number of sentences mostly contribute to the naturalness and readability of the translation.
  • Score 3 The translated text partially preserves the meaning, logic, and nuances of the original text. Some embedded clauses, conjunctions, and sentence connectors are translated inaccurately, resulting in awkward sentence structures in the target language. The number of sentences in the translation varies from the number of sentences in the original text without a clear demonstration of the enhancement in readability and naturalness.
  • Score 2 The translated text falls short of maintaining the meaning, logic, and nuances of the original text. The translation includes partial but critical mistranslation that makes it difficult to convey the meaning of the original. Many embedded clauses, conjunctions, and sentence connectors are translated inaccurately, which results in awkward and unnatural sentence structures in the target language. Variations in the number of sentences negatively affect the readability and comprehension of the translation.
  • Score 1 The translation completely fails to convey the meaning of the entire paragraph. The translated text demonstrates a significant misinterpretation of the original text due to a critical misunderstanding of the internal logic of its complex sentences. Embedded clauses, conjunctions, and sentence connectors are inaccurately translated, resulting in unnatural and awkward sentence structures in the target language. The translated text unnecessarily alters the number of sentences so that the overall comprehensibility and accuracy of the translation is critically compromised.
C.4 Content Accuracy
Preservation of the meaning of the text, with particular attention to the proper and comprehensive translation of various literary expressions from the original text. Proper translation might involve some additions and omissions, but it should refrain from adding or omitting substantial words or phrases that constitute the content and context of the original text.
  • Score 5 The translated text fully preserves the essential meaning of the original text, demonstrating thorough comprehension of the literary text. There may be additions or omissions of particular words in the translation, but these choices are made only when they significantly enhance content accuracy in translation. On the level of meaning and literary comprehension, none of the words or phrases in the original should be missing or misinterpreted in the translation.
  • Score 4 The translated text largely preserves the essential meaning of the original text, reflecting a good overall literary comprehension. There may be some minor omissions or additions of words or expressions, but these do not significantly alter the content and context of the original text.
  • Score 3 The translated text partially preserves the essential meaning of the original text, indicating some, if not critical, miscomprehension of the original text. There are some omissions or unnecessary additions of words or expressions, which somewhat alter the content and context of the original text.
  • Score 2 The translated text barely maintains the essential meaning of the original text due to omissions or inaccurate additions of words or phrases that reflect some literary misinterpretation. The translated text misinterprets some of words or idiomatic expressions in the original text, reflecting a significant misunderstanding of the literary work.
  • Score 1 The translated text fails to convey the essential meaning of the original text due to critical omissions or inaccurate additions, revealing a fundamental lack of literary comprehension of the original text. Meaningful words or phrases in the original are either missing or replaced with expressions that do not accurately capture the content and context of the original text.
Appendix D Detailed Rubric for Step 2
This is the rubric used for whether a given translation satisfied the criteria mentioned in LLM-generated questions.
Score 3: The translation meets the criteria. Even if the translation has some issues in other aspects including naturalness, it properly conveys the meaning and affective impact of the original text, especially with regards to the aspect emphasized in the question. This score does not indicate a perfect overall translation but success in addressing the specific aspect in the given question.
Score 2: The translation has some issues relevant to the specific aspect emphasized in the question, but mostly meets the criteria. The translation may include some awkward expressions when it comes to idiomatic use of Korean language, but mostly handles the focused aspect well with occasional mistakes or lack of precision.
Score 1: The translation fails to meet the criteria due to critical errors or mistakes relevant to the given criteria. The translation misses the mark especially with regards to the specific aspect emphasized in the given question due to a limited or flawed understanding of the particular element being evaluated.
Appendix E Annotation Guidelines and User Interface
E.1 Annotation Guidelines for Step 1: Common Mistake Patterns
Notice that the human annotators have access to both this annotator’s guideline and the rubric when they do annotation.
Basic Directions and Principles
  1. 1.
    You will be asked to evaluate literary translation on a 1-5 scale.
  2. 2.
    In the annotation UI, you will be given 1) the original text, 2) translation, 3) human translation(for reference), and 4) the criteria and rubric for evaluation.
  3. 3.
    You will be asked to give separate scores for all 4 criteria per translation. The rubric for each criterion will be shown on the screen. The human translation is provided on the UI just for reference, but note that they are not always scored as 5.
  4. 4.
    Scores for each assessment criterion should be independent of assessments in other criteria. (Even if the translation as a whole is not an ideal one, you should give score 4 or 5 if the translation aligns with the description for score 4 or 5 for a specific criterion.)
Detailed Guidelines and Examples
  1. 1.
    A lexical choice score of 4 is given when there are some awkward words (up to 3) in the translation, but the original meaning can be accurately captured. A lexical choice score of 3 is awarded when the translation contains a word (up to one) that misinterprets the original meaning but does not detract from the overall meaning.
  2. 2.
    Syntax and Grammar are separate from words and are scored based on whether the sentence form or structure is appropriate to capture meaning. For example, if the words are all translated correctly, but the sentence form and structure make it difficult to understand the meaning, a low score is given. In contrast, a high score is given if the sentence structure and grammar are appropriate, even if the words are not.
  3. 3.
    If there are no sentences in the paragraph that can be evaluated using the Honorifics criteria, then the score is 5.
  4. 4.
    If the use of honorifics is a translator’s choice, it can be given high scores even if it is not the same as human translation.
  5. 5.
    Content Accuracy is linked to the broader scope of accuracy and looks mainly at whether anything is missing/added in the translation. If nothing is missing/added, then evaluate how well the translation conveys the overall flow of the original, leaving out the details.
  6. 6.
    When evaluating lexical choices, phonetically writing an English word in Korean ({CJK*}UTF8mj팬트리>pantry) does not count as an error. However, it can be counted as an error if it hinders the reading experience because it is uncommon/awkward.
  7. 7.
    The honorifics in the narration were evaluated using the syntax criteria.
E.2 Annotation Guidelines for Step 2: LLM Generated Question Answering
Similar to Step 1 annotation, annotators also had access to the rubric alongside this guideline.
Basic Directions and Principles
  1. 1.
    You will be given the original text, a human translation of the given text, and a machine-translation of the given text.
  2. 2.
    You will then be given a question that addresses a specific aspect of the translation. Those specific aspects may include: historical and cultural context, imagery and descriptive quality, characterization, interpersonal dynamics, idiomatic naturalness, subtle implications, narrative pacing and rhythm, stylistic resonance, and coherence.
  3. 3.
    After reading the original text and the machine-translation carefully, give 1 to 3 scores based on your evaluation of the translation, focusing on the criteria relevant to the given question.
  4. 4.
    When you give scores, focus only on the specific aspect mentioned in the given question when you give scores.
  5. 5.
    Refer to the rubric above for what each score means.
Specific Guidelines
  1. 1.
    Score 3 does not necessarily signify “perfect translation.” Even when the translation includes some awkward expressions in Korean, give 3 if the translation satisfies the criteria conveyed through the given question.
  2. 2.
    When the given question does not seem to apply directly to the given passage, evaluate the overall quality of translation to give 1 to 3 scores.
  3. 3.
    When the question doesn’t seem to clearly reflect the content of the translation, make proper judgments about the assumptions of the question and answer accordingly.
  4. 4.
    If the question asks you to evaluate multiple facets, and part of the criteria is met, give 2
  5. 5.
    Interpersonal dynamics–consistent and proper=3 / consistent but not accurate, or proper but partly inconsistent=2 / inconsistent and exhibits misunderstanding=1
  6. 6.
    In terms of tone, if the overall flow matches the original, give 3.
  7. 7.
    Even when the translations are awkward in terms of idiomaticity, give 3 if it meets the criteria
  8. 8.
    When the translation totally fails to capture the content of the original text, give 1.
  9. 9.
    When there are multiple verbs in the translation and the question asks about using proper verb endings, consistent and proper=3, consistent but slightly awkward=2, inconsistent and/or awkward verb endings=1
  10. 10.
    When the translation does not include enough information to answer the questions (when the question is about a broader context that is not clear in the given paragraph) > evaluate the overall quality of the translation
  11. 11.
    In case there are random foreign words or expressions included in the translation: if the foreign language parts are limited to the names or other proper nouns, give 2 or 1 depending on the quality of the translation in terms of the aspect the question focuses on. Do not give 3 to these imperfect Korean translations. if one can understand the content and context of the overall translation (even if it displays awkwardness or some mistranslation), do not mark “the translation has serious issues.” Give it 2 or 1 depending on the quality of the translation especially in terms of the aspect the question focuses on. if the entire translation includes serious grammatical errors or complete mistranslation that makes it impossible to understand the entire passage, give 1.
  12. 12.
    If a significant portion of the translation is in languages other than Korean or if the text has missing parts (i.e., the entire last line/sentence is missing) mark “This translation has serious issues” and do not give 123 scores. Just click on Submit after checking “Yes” to “serious issues” question.
  13. 13.
    If there is something wrong with the question (i.e., the question is not a question but a category or a phrase that falls short of constituting a question), do not answer that question. If only one or two of the questions from the three questions have this issue, do not give 123 scores to those questions but give 123 scores to the rest of the questions. Do not mark the entire item as “This translation has serious issues.”
  14. 14.
    If the passage does not offer enough context to answer the questions properly (primarily because the question is about the broader context of the story), give 3 (unless the translation fails to convey the content of the original text).
  15. 15.
    Although it is difficult to ignore awkward expressions or grammatical imperfections in terms of fluency in Korean, always focus on the specific aspect the question asks you to evaluate.
E.3 Annotation Platform
We use Label Studio framework Tkachenko et al. (2020-2022), hosted on Huggingface platform777huggingface.co, to create an annotation platform for all our annotations.
E.4 Annotated Samples
After the annotation, and some post-processing to discard some samples, we had 188 data points for RULER (Step 1) and 543 data points for validating VERSE (Step 2).
We will release the human annotation for future researchers.
(a) Platform for Step 1 Annotation (1 to 5 MQM)
(b) Platform for Step 2 Annotation (1 to 3 MQM)
Figure 4: Annotation platform for annotations in Section 5
Appendix F Model Name Details
Here we list the full names of the models we used in the paper: gpt-4o-2024-05-13, claude-3-5-sonnet-20240620, claude-3-5-haiku-20241022, gpt-4o-mini-2024-07-18, claude-3-haiku-20240307, gemma-2-27b-it, gemma-2-9b-it, Llama-3.1-8B-Instruct.
F.1 Baseline Details
For the BERTScore, BLEURT, COMET baselines, we use the following versions used in the previous work Karpinska and Iyyer (2023):
BERTScore:
We used a roberta-large pretrained checkpoint from huggingface.
BLEURT:
We used bleurt-20 checkpoints from the official repository.
COMET:
wmt22-comet-da checkpoints.
Appendix G Prompting Methods Used for Generating Diverse Translations
To ensure that we have a diverse set of translations for the human evaluation stage mentioned in Section 5, we took several big and small language models and prompted them in various ways to obtain translations. We specifically used a combination following prompting techniques:
  1. 1.
    Sentence-wise prompting where we tokenize each paragraph into individual sentences and then give it to LLMs to translate.
  2. 2.
    Paragraph level prompting where we give LLMs the whole paragraph and ask it to translate in the target language.
  3. 3.
    We also incorporate zero shot and few shot prompting for the abovementioned scenario. Human inspection revealed that providing a few example translations always leads to better translation, as echoed in previous work like Zhang et al. (2023).
  4. 4.
    In a few variants we also add an LLM generated summary on top of few shot examples.
For the final analysis in Section 6 we generate translations using five-shot prompting with GPT4o summary appended. The five examples were sourced from stories other than the one that is being translated. We made sure to keep both dialogue-based and narrative-based paragraphs in the few shot examples.
All LLM translations were generated using the default temperature of the model (as opposed to 0.0 for evaluation).
Appendix H Additional Results and Ablations
H.1 LLM Generated Questions to Evaluate Literary Translations
H.1.1 Categories
These are the original human-curated categories of the LLM-generated questions:
"Historical and Cultural Context", "Imagery and Descriptive Quality", "Character Voice, Tone, and Individuality", "Interpersonal Communication and Hierarchy", "Linguistic and Idiomatic Naturalness", "Nuanced Interpretation including Subtle Implications", "Narrative Pacing and Rhythm", "Affective and Stylistic Resonance", "Overall Consistency and Cohesion"
The categories have been abridged on all the plots for the sake of visualization.
H.1.2 Quality of the LLM Generated Questions
On average, GPT4-o generates 9.6$\pm$3.4 questions per paragraph. A careful examination by a human expert on 50 randomly selected questions finds 44 questions to be “very helpful" to evaluate literary quality. The six remaining questions, while still relevant, are deemed either “too generic" (4) or unanswerable based on the context of that paragraph alone (2). For example, GPT4-o generates multiple questions about the historical and cultural context for a story set in the mid-19th century America ("The Passing of Grandison"), whereas it generates questions about nuanced interpretations of certain phrases or the use of specific imagery for a story with rich poetic descriptions ("The Ones Who Walk Away from Omelas").
H.1.3 Zero Shot Classification Accuracy
Figure 5: Confusion matrix of the zero-shot question classification. The category names are slightly abridged to fit in the figure. Notice that most apparent misclassifications are usually in a very closely related class and human inspection revealed that in most cases the misclassified class can also be considered valid.
We use GPT4-o to classify the questions. We passed all the questions belonging to a particular paragraph to the LLM and got them classified together to save token cost. Note that, in a few cases ( 5%) the pipeline failed to generate questions that fit the formatting requirement for the classification pipeline and we discarded those questions from our analysis. In most of these cases, the question generator model generated something that couldn’t be parsed properly. This is, however, not a fundamentally difficult issue to solve.
Figure 5 shows the confusion matrix of GPT4o’s ability to classify questions. Here the average accuracy was 0.77 with precision = 0.76, recall = 0.78 and f1 = 0.75.
H.2 Detailed Analysis of each criterion of Step 1
We observe some interesting patterns across the score distribution for Step 1 RULER, especially for Korean honorifics.
Figure 6: The confusion matrices show the scores given by two human annotators and LLM for Korean Honorific. As the upper row shows there is good agreement between the annotators that some paragraphs have very poor quality in terms of honorifics (notice the dark blue corner on the upper left). However, as the second row shows, GPT4-o fails to capture this (the upper left corner is lighter). The matrices are normalized column-wise (axis=0)
The Case of Korean Honorific
Honorific is one of the most important aspects of Korean translation, however, honorifics are only relevant in dialogue-like texts. If the text does not involve dialogue, the honorific score is meant to be given 5 according to our rubric. This means that there were a large number of evaluation data points where honorific was not important.
In Korean, the reading experience is heavily affected if proper honorific is not used, it is crucial for the translation judge to be able to capture those mistakes. However, the best models’ precision and recall for low honorific labels is very low (see the bottom row of Figure 6). Table 3 shows aggregated F1 values for all the criteria for each label and GPT4-o has poor F1 values for scores 1, 2, and 3.
Figure 7: Each row represents an agreement between annotator 1 and 2 and comparison with annotator 2’s agreement (left) with GPT4o for each criterion (right). Notice that GPT4-o is better at matching the diagonal patterns in human agreement for these criteria.
Notice that in Figure 7, the language model is better at matching the preference distribution of humans (which is almost diagonal) in other dimensions.
H.3 Details for of VERSE (Step 2)
Table 8 shows classification statistics averaged over three annotators for each score class. Models tend to do worse when humans gave a Score 1, that is, models struggle to spot a bad translation. This is similar to the behavior observed for honorifics in RULER (Step 1).
Human-Human Human-Model
Label Support Prec Rec F1 Prec Rec F1
1 197/189 0.81 0.75 0.78 0.69 0.06 0.11
2 182/171 0.52 0.46 0.49 0.32 0.78 0.45
3 185/204 0.64 0.77 0.70 0.68 0.45 0.54
Avg 188/188 0.66 0.66 0.66 0.56 0.43 0.42
Table 8: Comparison of classification metrics between human annotators (Human-Human) and GPT4-o no CoT predictions (Human-Model). Support shows the average number of instances for human/model evaluations.
H.4 Ablations for RULER (Step 1)
CoT hurts across all models
Table 9 of Appendix H, demonstrates chain of thought prompting seems to hurt performance across all models, except for some models which see a minor improvement in honorifics when used CoT. This observation is at odds with the common practice of LLM-as-a-judge literature and deserves further study.
Honorifics Lexical Syntax Content Aggregated Avg.
Model $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE Avg. $\tau$ Avg. $\rho$ Avg. MSE
gpt-4o .50 .55 2.10 .68 .77 0.71 .62 .71 0.95 .70 .79 0.81 .62 .70 1.14
w/ CoT .44 .48 2.00 .62 .71 0.86 .59 .68 1.02 .61 .70 1.12 .56 .64 1.25
claude-3-5-sonnet .39 .41 2.30 .58 .65 1.17 .57 .65 1.20 .57 .64 1.38 .53 .59 1.51
w/ CoT .45 .48 2.15 .53 .61 1.21 .55 .63 1.51 .56 .63 1.58 .52 .59 1.61
claude-3-5-haiku .41 .44 1.86 .54 .60 1.37 .57 .64 1.23 .55 .62 1.42 .52 .58 1.47
w/ CoT .44 .47 1.92 .53 .60 1.24 .46 .53 1.62 .32 .36 1.91 .44 .49 1.67
gpt-4o-mini .26 .29 2.84 .62 .70 0.93 .60 .68 0.91 .60 .67 1.10 .52 .58 1.44
w/ CoT .28 .31 3.36 .46 .51 1.51 .47 .54 1.36 .60 .69 1.00 .45 .51 1.81
claude-3-haiku .20 .21 2.24 .45 .50 1.57 .42 .47 1.47 .50 .56 1.55 .39 .43 1.71
w/ CoT .16 .17 2.28 .40 .44 1.74 .44 .49 1.61 .39 .43 1.89 .34 .38 1.88
gemma-2-27b .33 .36 2.16 .46 .51 1.27 .48 .54 1.24 .48 .54 1.26 .44 .49 1.48
w/ CoT .28 .31 2.11 .37 .42 1.46 .35 .40 1.45 .52 .58 1.34 .38 .43 1.59
gemma-2-9b .27 .28 2.44 .42 .47 1.41 .35 .40 1.48 .52 .58 1.35 .39 .43 1.67
w/ CoT .32 .35 2.01 .44 .49 1.35 .29 .33 1.65 .31 .35 1.59 .34 .38 1.65
llama-3-8b -.02 -.02 2.36 .37 .41 1.58 .28 .32 1.67 .21 .23 2.09 .21 .23 1.92
w/ CoT .12 .13 2.50 .13 .14 1.90 .20 .23 1.82 .26 .29 1.84 .18 .20 2.01
Table 9: Performance comparison across different linguistic aspects with and without Chain-of-Thought (CoT) prompting. Gray rows indicate results with CoT prompting. $\tau$: Kendall’s Tau, $\rho$: Spearman’s correlation, MSE: Mean Squared Error.
Specific Rubric is Important
As shown in Table 10 providing fine-grained human-curated rubric is important for achieving high correlation with GPT4-o. If the detailed rubric is removed. However, different models are affected differently and for some models having a rubric is critical.
Honorifics Lexical Syntax Content Aggregated Avg.
Model $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE Avg. $\tau$ Avg. $\rho$ Avg. MSE
gpt-4o .50 .55 2.10 .68 .77 0.71 .62 .71 0.95 .70 .79 0.81 .62 .70 1.14
w/o rubric .22 .26 4.67 .68 .76 0.81 .62 .71 0.88 .65 .73 0.85 .54 .61 1.80
w/o human TR .50 .53 2.04 .68 .77 0.82 .65 .75 1.10 .69 .77 1.01 .63 .71 1.24
gpt-4o-mini .26 .29 2.84 .62 .70 0.93 .60 .68 0.91 .60 .67 1.10 .52 .58 1.44
w/o rubric .20 .22 3.36 .63 .70 0.92 .62 .69 0.93 .62 .69 0.96 .52 .58 1.54
w/o human TR .30 .34 2.64 .61 .68 0.94 .58 .66 0.99 .60 .68 1.17 .52 .59 1.43
gemma-2-27b-it .33 .36 2.16 .46 .51 1.27 .48 .54 1.24 .48 .54 1.26 .44 .49 1.48
w/o rubric .31 .33 3.96 .51 .57 1.50 .53 .60 1.17 .58 .66 1.10 .48 .54 1.93
w/o human TR .37 .40 1.96 .52 .58 1.12 .52 .59 1.09 .54 .62 1.17 .49 .55 1.34
gemma-2-9b-it .27 .28 2.44 .42 .47 1.41 .35 .40 1.48 .52 .58 1.35 .39 .43 1.67
w/o rubric .20 .21 3.87 .27 .30 1.81 .26 .29 1.72 .20 .22 1.77 .23 .25 2.29
w/o human TR .29 .32 2.15 .54 .60 1.37 .49 .54 1.53 .51 .57 1.54 .46 .51 1.65
llama-3-8b-it -.02 -.02 2.36 .37 .41 1.58 .28 .32 1.67 .21 .23 2.09 .21 .23 1.92
w/o rubric -.04 -.05 3.21 .17 .19 2.03 .13 .15 2.19 .15 .17 2.23 .10 .11 2.42
w/o human TR .11 .12 2.39 .36 .40 1.65 .31 .35 1.74 .15 .17 2.23 .23 .26 2.00
Table 10: (Step 1) Ablation study results comparing model performance under different conditions. Base model performance is shown in white rows, while ablation conditions (without rubric and without human training) are shown in gray rows. Note that for most models, removing the rubric has a bigger effect than removing the human translation
Human Translation is Not Necessary But Good to Have
As seen in Table 10, while human rubrics are good to have, removing them doesn’t lead to a significant performance drop (in terms of correlation) in most criteria. In fact, some models do slightly better without human translation. This is good news since having a human translation for literary text is not possible for most of cases. This might also open the door to evaluating different versions of human translations and quantitatively evaluate their differences. However, note that Avg. MSE is higher for w/o human reference cases, which usually means LLMs are biased towards giving higher scores to machine translation when human translation is not provided.
Honorifics Lexical Syntax Content Aggr. Avg.
Model $k$ $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE Avg. $\tau$ Avg. $\rho$ Avg. MSE
gpt-4o 0 .50 .55 2.10 .68 .77 0.71 .62 .71 0.95 .70 .79 0.81 .62 .70 1.14
5 .49 .52 1.98 .68 .77 0.75 .63 .72 0.96 .67 .76 0.79 .62 .69 1.12
10 .49 .53 2.00 .69 .78 0.75 .64 .73 0.97 .65 .73 0.88 .62 .69 1.15
15 .48 .52 2.15 .68 .78 0.75 .64 .74 1.02 .66 .75 0.85 .62 .70 1.19
20 .44 .48 2.48 .67 .76 0.80 .64 .73 0.95 .66 .75 0.89 .60 .68 1.28
gpt-4o-mini 0 .26 .29 2.84 .62 .70 0.93 .60 .68 0.91 .60 .67 1.10 .52 .58 1.44
5 .30 .33 3.40 .54 .61 1.19 .60 .67 0.93 .59 .66 1.02 .51 .57 1.64
10 .30 .33 3.31 .52 .59 1.31 .53 .60 1.10 .59 .66 1.01 .49 .55 1.68
15 .33 .37 3.14 .48 .54 1.36 .50 .57 1.17 .64 .71 0.92 .49 .55 1.65
20 .25 .27 3.16 .51 .57 1.36 .51 .58 1.14 .59 .66 1.01 .46 .52 1.67
Llama-3.1-8B-It 0 -.02 -.02 2.36 .37 .41 1.58 .28 .32 1.67 .21 .23 2.09 .21 .23 1.92
5 .19 .20 2.32 .20 .23 1.97 .12 .14 2.21 .23 .26 2.03 .19 .21 2.13
10 .10 .10 2.25 .13 .14 2.05 .14 .16 2.22 .14 .16 2.17 .13 .14 2.17
15 .03 .03 2.32 .18 .20 2.01 .16 .18 2.19 .14 .16 2.19 .13 .14 2.18
20 .24 .26 2.22 .24 .27 1.89 .19 .22 2.08 .23 .26 2.04 .23 .25 2.06
Table 11: (Step 1) Ablation showing the effect of varying $k$ (number of demonstrations) on model performance across different metrics (no chain of thought). Notice that for large models, providing examples doesn’t make a significant difference. For the small Llama-3.1 model, the pattern is more inconsistent, but no significant gain was observed either. $\tau$: Kendall’s Tau, $\rho$: Spearman’s correlation, MSE: Mean Squared Error. Note that gemma models were excluded from this analysis because of their shorter context windows. Human translations were evaluated in a no-reference format.
Providing few shot examples doesn’t help
We experimented with few shot prompting by providing a few examples (without CoT). However, as the results in Table 11 show there is no significant performance gain across models when few annotated examples were given. In fact, some models tend to do better without few shot examples. We kept the few shot examples the same across all models and made sure that there was at least one example from each score (1 to 5) among the examples.
H.5 Ablations for VERSE (Step 2)
Model $\tau$ $\rho$ MSE
gpt-4o 0.49 0.54 0.61
w/ CoT 0.45 0.50 0.69
claude-3-5-sonnet 0.51 0.54 0.73
w/ CoT 0.36 0.39 0.65
claude-3-5-haiku 0.37 0.40 0.77
w/ CoT 0.39 0.42 0.67
gpt-4o-mini 0.37 0.40 0.61
w/ CoT 0.34 0.36 0.63
claude-3-haiku 0.20 0.21 0.82
w/ CoT 0.17 0.19 0.89
gemma-2-27b 0.13 0.14 0.71
w/ CoT 0.09 0.10 0.78
gemma-2-9b 0.09 0.09 0.71
w/ CoT 0.16 0.17 0.88
llama-3-8b 0.05 0.05 1.08
w/ CoT 0.14 0.16 0.90
Table 12: (Step 2) Performance of different models with and without Chain-of-Thought (CoT) prompting. The effect of CoT is inconsistent and depends on the specific models. Some models do benefit from CoT.
CoT only helps some models
As seen in Table 12, chain of thought reasoning doesn’t GPT series models much, in fact, it hurts. However, for smaller models, doing chain of thought boosts the performance from completely random to 0.20.2~{}0.20.2. However, even for the models that benefit from CoT (in terms of correlation), some of them have higher MSE, which usually means they tend to give higher scores to machine translations.
Default No Rubric No Human Tr
Model $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE $\tau$ $\rho$ MSE
gpt-4o 0.49 0.54 0.61 0.08 0.09 1.22 0.52 0.57 0.59
gpt-4o-mini 0.37 0.40 0.61 0.02 0.02 1.05 0.42 0.46 0.60
gemma-2-27b 0.13 0.14 0.71 0.07 0.08 0.70 0.25 0.26 0.75
gemma-2-9b 0.09 0.09 0.71 0.02 0.02 0.79 0.17 0.18 0.75
llama-3-8b 0.05 0.05 1.08 -0.02 -0.02 0.73 0.06 0.06 1.08
Table 13: (Step 2) Performance comparison across different models and configurations. No Rubric refers to when human written rubric was not provided and No Human Tr means human reference translation was not provided. Notice that while having access to human translation is not critical for large models, having access to human-curated rubic is essential for all models.
Extensive rubric is critical but reference translation is not
As seen in Table 13, not providing human written criteria leads the correlation to drop to zero. A closer inspection revealed that when a specific grading rubric is not provided, the model gives most of the translations a 3 out of 3. While the number of 2s doesn’t change much, not providing specific rubrics decreases the number of 1s the model provides.
The table also shows that not providing human translation works better than providing reference translation.
Model k $\tau$ $\rho$ MSE
gpt-4o 0 0.49 0.54 0.61
5 0.47 0.51 0.65
10 0.52 0.57 0.58
15 0.48 0.53 0.61
20 0.50 0.54 0.63
gpt-4o-mini 0 0.37 0.40 0.61
5 0.20 0.21 0.75
10 0.14 0.15 0.74
15 0.20 0.22 0.73
20 0.19 0.21 0.77
llama-3-8b-it 0 0.05 0.05 1.08
5 0.04 0.04 1.59
10 0.06 0.07 1.31
15 0.06 0.06 0.82
20 0.05 0.05 0.77
Table 14: (Step 2) Performance comparison across k shot examples provided (in no COT setting) in answering LLM generated evaluation questions. Note that gemma models are excluded because of their shorter context length.
Providing few shot examples doesn’t make a big difference
As Table 14 shows, providing a few shot examples doesn’t significantly boost performance across models.
Appendix I Additional Results from Discussion (Section 6)
I.1 Translation Generation
We selected the bottom 10 stories from Table 7 and generated their translations using language models and Google Translate. We use the top 5 stories of Table 7 to obtain few shot examples. We used one paragraph from each story for few shot prompting. One example paragraph translated across different models can be found in Table 17.
Figure 8: Fine-grained evaluation of 8 stories using VERSE. Notice that models score widely differently for different stories. The difference is more pronounced for less powerful model, for whom evaluator LLM is more sensitive. $n$ refers to the number of questions belonging to that category. The legend on the bottom left shows which color belongs to which model.
I.2 LLM’s tendency to overestimate scores
As shown in Table 15 and 16, evaluator LLM GPT4-o tends to overestimate the quality of LLM generated translations over human translations. An interesting pattern is observed with the direction of overestimation. For RULER (Step 1) (Table 15) scores, providing reference translation decreases the average score, while for LLM generated QnA in Step2 (Table 16), the average score increases. This is in line with the results observed in the ablations in Table 10 and 13, where MSE increases when reference is not provided for RULER (Step 1), and decreases for VERSE (Step 2).
Model Hon. Syn. Lex. Con.
gpt-4o 99.49 97.44 94.36 97.95
w/ Ref 98.97 86.67 82.56 90.77
claude-3.5-sonnet 100.00 98.97 98.97 98.97
w/ Ref 99.49 95.90 90.77 96.92
claude-3.5-haiku 100.00 94.87 96.41 96.92
w/ Ref 98.97 87.69 82.56 91.28
claude-3-haiku 98.46 88.72 85.64 89.23
w/ Ref 96.92 80.51 76.92 81.54
gpt-4-mini 99.49 90.26 88.21 92.31
w/ Ref 98.97 80.51 75.90 86.15
gpt-3.5 91.79 57.44 56.92 60.00
w/ Ref 88.21 51.79 51.28 54.87
google-translate 97.95 71.79 69.23 76.41
w/ Ref 95.90 66.15 63.59 68.72
human 98.97 88.21 89.74 83.08
Table 15: Score estimation averaged over all 10 stories (in the percentage) of individual models with and without providing reference prompting across different metrics: honorifics (Hon.), syntax (Syn.), lexical choices (Lex.), and content preservation (Con.). Gray rows indicate the scores computed when reference translation was also provided to $\mathrm{LM}_{E}$. Note that providing the translation decreases the absolute score. The human translation was evaluated in a reference-free setting.
Model Style Char. Hist. Img. Comm. Ling. Pace Nuance Cons. Mean
gpt-4o 89.70 92.16 90.58 90.76 93.09 88.36 96.42 92.00 98.64 92.41
w/ Ref 93.21 94.68 90.12 92.28 94.87 90.42 98.02 94.04 99.42 94.12
claude-3.5-sonnet 94.42 95.35 92.51 94.12 96.12 93.21 98.91 95.29 100.00 95.55
w/ Ref 95.88 96.87 92.28 95.07 96.64 94.42 99.43 96.71 100.00 96.37
claude-3.5-haiku 90.30 92.46 90.47 91.70 93.53 87.39 97.13 92.86 99.03 92.76
w/ Ref 94.18 94.93 90.00 93.36 94.87 89.58 98.47 94.43 99.61 94.38
claude-3-haiku 85.82 87.06 86.84 85.33 89.40 80.61 93.93 87.45 96.51 88.11
w/ Ref 88.48 89.67 85.38 86.68 90.55 83.52 94.83 89.10 97.29 89.50
gpt-4-mini 82.79 87.15 85.91 85.73 88.06 81.82 92.02 87.37 96.71 87.51
w/ Ref 87.15 90.09 85.44 87.21 90.70 84.12 94.70 89.73 98.64 89.75
gpt-3.5 64.48 67.03 67.54 65.68 68.06 63.27 69.28 67.29 75.39 67.56
w/ Ref 64.61 66.70 64.91 63.53 67.91 62.55 68.97 67.22 72.87 66.58
google-translate 71.39 73.14 74.68 74.79 76.21 68.00 79.18 75.76 86.63 75.53
w/ Ref 72.24 74.42 73.04 74.07 75.78 67.64 78.86 74.90 84.88 75.09
human 88.48 89.40 87.02 88.29 90.84 84.12 94.06 88.78 95.54 89.62
Table 16: (Step 2) Aggregated scores for the ten stories on Style, Character Voice (Char.), Historical Context (Hist.), Imagery (Img.), Communication (Comm.), Linguistic accuracy (Ling.), Pacing (Pace), Nuanced Interpretation (Nuance), and Consistency (Cons.). Gray rows indicate aggregated scores when the reference human translation was also provided. Scores for human translation were calculated in reference free manner.
Notice, however, that the relative rank of the models changes in a few cases in these two settings, especially in the middle ranks. However, the best and worst-performing models’ ranks do not change whether reference translation is provided or not.
This observation supports our previous claim in Section 6 that GPT4-o as a judge can differentiate between obviously good and obviously bad translations, however, its capability to detect differences in similarly (high) quality translations is not as robust.
I.3 The Curious Case of CoT
While Chain of Thought (CoT) seems to help a range of reasoning problems and is a standard in LLM as a judge setting, we find CoTs not too helpful for our task, in fact in some cases doing CoT actually hurts performance. This might be related to some recent work that suggests CoT only helps mathematical and symbolic reasoning tasks Sprague et al. (2024) and might even hurt in some cases Liu et al. (2024). In this section, we select some of the GPT4-o intermediate chain of thought reasoning and conduct a qualitative analysis.
LLMs’ analyses of the atmosphere, context, and character relationships in the original texts prove to be quite accurate; however, in their evaluation of Korean translations based on those criteria, they often fail to make nuanced decisions, especially when it comes to judging whether certain expressions sound natural in colloquial Korean. On the level of lexical choice, LLMs demonstrate a superb understanding of the nuances of Korean words and are adept at assigning culturally and historically appropriate titles to character names. In their evaluations, LLMs often employ vague phrases such as “confusing and not idiomatic” for clearly ungrammatical sentences or label certain phrases or sentences as “accurately translated” even when the translations sound critically awkward in Korean. CoTs provided by LLMs also demonstrate a theoretical understanding of when different verb endings should be used for proper levels of politeness and formality—they recognize "{CJK*}UTF8mj하오", "{CJK*}UTF8mj없소" as older, formal language and mark "{CJK*}UTF8mj열어" or "{CJK*}UTF8mj죽고 말 거야" as more contemporary informal language—, but they turn out to be relatively incapable of making proper judgments about when and how to use each level of formality appropriately.
Appendix J Prompts Used for the Experiments
J.1 Prompt Used to Generate Evaluation Questions
You are supposed to evaluate a literary translation that has already been done from English to Korean.
This is a summary of the original story in English:
<SUMMARY>
<An LLM generated story summary is provided>
</SUMMARY>
Your job is to evaluate whether a particular paragraph in the story is *successfully* translated or not.
But before you see the translation, write down the criteria *specific to this paragraph* that need to be met for the translation to be successful in Korean.
Your criteria should focus on these things:
  1. 1.
    Characterization of individual characters: This category has to do with physical description, psychological depth, and the developments each individual character goes through in the given passage. Subtle nuances and uniqueness of the characters captured in the original text should be properly translated.
  2. 2.
    Interpersonal Dynamics: This category focuses on capturing the level of tension, affection, familiarity, or any other power dynamics at work in the interactions (including dialogues) between characters. Special attention needs to be paid to the social and cultural roles played by the characters in the story and the nature of the relationship depicted in the story.
  3. 3.
    Narrative Voice: This category pays particular attention to the tone of the story that distinguishes this story from any other stories. The narrative voice is largely determined by the verb endings, and the verb endings should properly distinguish between formal, informal, colloquial, poetic, or regional voices used in each work.
  4. 4.
    Unique Use of Expressive Language: This category deals with the author’s unique, literary use of language that distinguishes their work from other writers’ work. The story’s use of figurative language or the quality of descriptions in the story largely determines how and in what ways the story is unique, and such uniqueness should be properly captured in a successful translation. Word choices are particularly important here because.
  5. 5.
    Understanding the Atmosphere of the Setting: This category examines how each translation takes into account the historically and culturally specific elements from the text. If the story has a unique and distinctive setting that greatly contributes to the literary impact of the story, the historical and cultural context of the setting needs to be considered as one of the most important aspects of a work that is to be translated.
  6. 6.
    Appropriate Portrayal of the Event (Plot): First of all, what happens in the story should be translated accurately without any misinterpretation or miscomprehension, but the pacing, as well as the mood of the event, too, need to be taken into account when evaluating quality of translations. If the given passage includes subtle depictions of a complex event depicted in a unique way, a successful translation should pay particular attention to accurate portrayal of the event.
Note that not all paragraphs will have all of these elements. Some paragraphs may have one and some may have all elements. You can also go beyond these categories if you think there are other important elements that need to be considered. The above is just for reference. Use the best of your judgment to make sure which elements are important for the *given paragraph* for the given story
The given paragraph is:
<PARAGRAPH>
<The original paragraph>
</PARAGRAPH>
Be specific to the story and the target language.
DO NOT mention about consulting other people or resources. YOU are the expert here. Don’t mention about footnotes or explanations. ONLY return the criteria, and nothing else.
J.2 Prompts Used in Classifying Questions
You are given a list of questions that are used to evaluate literary translations.
Your job is to classify those questions into one of these categories:
categories = ["Historical and Cultural Context", "Imagery and Descriptive Quality", "Character Voice, Tone, and Individuality", "Interpersonal Communication and Hierarchy", "Linguistic and Idiomatic Naturalness", "Nuanced Interpretation including Subtle Implications", "Narrative Pacing and Rhythm", "Affective and Stylistic Resonance", "Overall Consistency and Cohesion" ]
One question belongs to ONLY one category.
The question list is:
<Question List>
<questions for each paragraph given as a python list format>
</Question List>
Return the category for each question as a list. DO NOT return anything else. ONLY the list. DO NOT return triple quotes. ONLY the list.
J.3 Prompt for Step 1 MQM (1 to 5) Grading
This is the prompt for the experiment results for Table 4. Note that ablations use modified versions of this prompt.
Prompt:
You are supposed to evaluate literary translations that has already been done from Korean to English.
You have to rate it between 1 to 5 according to a grading criteria.
This is the grading criteria for the translation: <CRITERIA>
<The detailed criteria goes either for Honorifics, Lexical Choice, Syntax, or Content Accuracy>
</CRITERIA>
Now, this is the translation you will actually evaluate:
<ORIGINAL>
<Original Paragraph>
</ORIGINAL>
<TRANSLATION>
<Translated Paragraph>
</TRANSLATION>
<HUMAN TRANSLATION>
<Human Translation Provided>
</HUMAN TRANSLATION>
You have to evaluate the given translation and NOT the human translation. The human translation is provided here just for reference, but note that they are not always scored as 5.
Choose among <1>, <2>, <3>, <4> or <5> as your final answer according to the rubric.
JUST RETURN THE FINAL ANSWER. NOTHING ELSE. NO OTHER TEXT. JUST THE FINAL ANSWER inside <>
Put your final answer within <>.
Appendix K Examples of LLM Generated Questions in VERSE (Step 2)
Paragraph
(The Red Masque of Death)
The “Red Death” had long devastated the country. No pestilence had been ever so fatal, or so hideous. Blood was its Avator and its seal — the redness and the horror of blood. There were sharp pains, and sudden dizziness, and then profuse bleedings at the pores, with dissolution. The scarlet stains upon the body and especially upon the face of the victim, were the pest-ban which shut him out from the aid and from the sympathy of his fellow-men. And the whole seizure, progress and termination of the disease were the incidents of half an hour.
Generated Questions:
  • Does the translation effectively personify the "Red Death" with its fatal and hideous qualities, and depict its impacts on the victims, particularly the scarlet stains and their social ostracization?
  • Does the translation clearly convey the societal response to the victims, reflecting the horror and social consequences of the plague?
  • Does the translation capture the formal, ominous tone of the original text, using verb endings that convey a sense of dread and formality?
  • Does the translation accurately render the vivid and figurative descriptions, such as "Blood was its Avator and its seal," and "scarlet stains," preserving the intensity of the original imagery?
  • Does the translation successfully convey the bleak, terrifying atmosphere of a country ravaged by an unprecedented and gruesome plague, with a palpable sense of pervasive dread and isolation?
  • Does the translation accurately portray the sequence of events—sharp pains, sudden dizziness, profuse bleeding, and swift death within half an hour— while maintaining the pacing and horror of the disease’s rapid progression?
Paragraph
(Hell-Heaven)
Pranab Chakraborty wasn’t technically my father’s younger Pranab brother. He was a fellow Bengali from Calcutta who had washed up on the barren shores of my parents’ social life in the early seventies, when they lived in a rented apartment in Central Square and could number their acquaintances on one hand. But I had no real uncles in America, and so I was taught to call him Pranab Kaku. Accordingly, he called my father Shyamal Da, always addressing him in the polite form, and he called my mother Boudi, which is how Bengalis are supposed to address an older brother’s wife, instead of using her first name, Aparna. After Pranab Kaku was befriended by my parents, he confessed that on the day we met him he had followed my mother and me for the better part of an afternoon around the streets of Cam- bridge, where she and I tended to roam after I got out of school. He had trailed behind us along Massachusetts Avenue and in and out of the Harvard Coop, where my mother liked to look at discounted housewares. He wandered with us into Har- vard Yard, where my mother often sat on the grass on pleasant days and watched the stream of students and professors filing busily along the paths, until, finally, as we were climbing the steps to Widener Library so that I could use the bathroom, he tapped my mother on the shoulder and inquired, in English, if she might be a Bengali. The answer to his question was clear, given that my mother was wearing the red and white bangles unique to Bengali married women, and a common Tangail sari, and had a thick stem of vermilion powder in the center parting of her hair, and the full round face and large dark eyes that are so typical of Bengali women. He noticed the two or three safety pins she wore fastened to the thin gold bangles that were behind the red and white ones, which she would use to replace a missing hook on a blouse or to draw a string through a petticoat at a moment’s notice, a practice he associated strictly with his mother and sisters and aunts in Calcutta. Moreover, Pranab Kaku had overheard my mother speaking to me in Ben- gali, telling me that I couldn’t buy an issue of Archie at the Coop. But back then, he also confessed, he was so new to America that he took nothing for granted and doubted even the obvious.
Generated Questions:
  • Does the translation accurately depict Pranab’s role and background?
  • Does the translation properly represent the unique characteristics of the narrator’s family?
  • Does the translation include a detailed description of the mother’s physical appearance and attire, emphasizing her Bengali identity?
  • Does the translation correctly use familial and respectful titles such as "Pranab Kaku," "Shyamal Da," and "Boudi"?
  • Does the translation convey the respect and formality in Pranab’s interactions with the narrator’s parents?
  • Does the translation maintain the tone and perspective of the original narrator?
  • Are the verb endings used in the translation appropriate to reflect the narrator’s voice, capturing both the formality and familial warmth present in the original text?
  • Does the translation preserve the descriptive language and cultural references such as "red and white bangles," "Tangail sari," and "vermilion powder"?
  • Does the translation accurately convey specific details like the practice with safety pins and Pranab’s doubts, highlighting his recent immigration experience?
  • Does the translation clearly depict the early seventies’ immigrant life and the setting of Central Square?
  • Does the translation accurately render the cultural and social isolation of the Bengali immigrants in that era, reflected in their limited social circle?
  • Does the translation precisely portray the sequence of events, especially how Pranab followed the narrator and her mother?
  • Does the translation maintain the suspense and curiosity of Pranab leading up to his inquiry about the mother being Bengali?
Paragraph
(The Garden Party)
And after all the weather was ideal. They could not have had a more perfect day for a garden-party if they had ordered it. Windless, warm, the sky without a cloud. Only the blue was veiled with a haze of light gold, as it is sometimes in early summer. The gardener had been up since dawn, mowing the lawns and sweeping them, until the grass and the dark flat rosettes where the daisy plants had been seemed to shine. As for the roses, you could not help feeling they understood that roses are the only flowers that impress people at garden-parties; the only flowers that everybody is certain of knowing. Hundreds, yes, literally hundreds, had come out in a single night; the green bushes bowed down as though they had been visited by archangels.
Generated Questions:
  • Does the translation maintain the slightly formal, reflective narrative tone appropriate for a literary description of the setting?
  • Does the translation capture the unique descriptive style of the original, including personification and vivid imagery?
  • Are phrases like "the sky without a cloud," "the blue was veiled with a haze of light gold," and "the green bushes bowed down as though they had been visited by archangels" translated in a way that maintains their evocative impact?
  • Does the translation effectively reflect the idyllic, picturesque garden-party setting described in the original text?
  • Does the translation convey the serene and perfect weather, the meticulous care of the gardener, and the vivid, almost supernatural bloom of the roses?
  • Does the translation ensure the reader feels the anticipation and effort that has gone into creating a perfect day?
  • Are the word choices and phrases in the translation reflective of the lush, detailed imagery of the original text, preserving the literary quality and nuanced meaning?
Paragraph
(The Passing of Grandison)
"I confess," he admitted, "that while my principles were against the prisoner, my sympathies were on his side. It appeared that he was of good family, and that he had an old father and mother, respectable people, dependent upon him for support and comfort in their declining years. He had been led into the matter by pity for a negro whose master ought to have been run out of the county long ago for abusing his slaves. If it had been merely a question of old Sam Briggs’s negro, nobody would have cared anything about it. But father and the rest of them stood on the principle of the thing, and told the judge so, and the fellow was sentenced to three years in the penitentiary."
Generated Questions:
  • Does the translation effectively convey the speaker’s internal conflict between his principles and his sympathies?
  • Are the perceptions of the prisoner, his family, and old Sam Briggs accurately depicted?
  • Is the tension between the speaker’s principles and his sympathy towards the prisoner clearly conveyed?
  • Are the power dynamics between the judge, the father, and the rest of the community accurately translated?
  • Is the confessional yet formal tone of the narrative voice maintained in the translation?
  • Is the speaker’s regret and understanding of the harshness evident?
  • Does the translation preserve the figurative or descriptive language that highlights the speaker’s conflicted feelings and the societal norms being questioned?
  • Are the word choices in the translation reflective of the gravity of the confession and the emotional undertones?
  • Does the translation reflect the historical and cultural context of the mid-19th century, particularly regarding attitudes toward slavery and justice?
  • Are the social and legal norms of the time evident in the translation?
  • Are the factual elements of the confession, including the prisoner’s background, the abuse of slaves, and the ultimate sentencing, accurately translated?
  • Is the pacing and mood of the event, including the community’s stance and the judge’s decision, preserved in the translation?
Appendix L Qualitative Examples
Table 17: Examples of translations made by different models from Section 6. This example is from the Tower of Babylon story. Human translation is provided at the bottom.
Table 18: Chain of Thought Example on Honorific criteria of Step 1. Example is taken from A Retrieved Reformation Story.
Table 19: Chain of Thought Example on Lexical Choices criteria of Step 1. Example is taken from A Retrieved Reformation Story.
Table 20: Chain of Thought Example on Syntax criteria of Step 1. Example is taken from A Retrieved Reformation Story.
Table 21: Chain of Thought Example on Content Accuracy criteria of Step 1. Example is taken from A Retrieved Reformation Story.
Table 22: Chain of Thought Example on a LLM Generated Question of Step 2. Example is taken from A Retrieved Reformation Story.
Table 23: Chain of Thought Example on a LLM Generated Question of Step 2. Example is taken from A Retrieved Reformation Story.
Table 24: Chain of Thought Example on a LLM Generated Question of Step 2. Example is taken from A Retrieved Reformation Story.